# Enhancing Cryptocurrency Sentiment Analysis with Multimodal Features

### v1

@@ Abstract
[inst2]organization=Department of Science, Queensland University of Technology, city=Brisbane, state=Queensland, country=Australia
@@
Research highlight 1
Research highlight 2
@@ 3.1 Data Collection
Account Name Number of Followers
crypto.lucho 2.7M
CryptoMasun 1.4M
CryptoProjects 1.3M
thecryptonetwork 1.0M
cryptosanti 914K
cryptorobles 886K
suppomancrypto 344K
Virtual Bacon 338K
thecryptohippie 288K
Cryptowendyo 277K
@@
For each group, one representative coin was chosen to conduct dependency analysis between sentiment signals and forecast modeling, providing insights into the differing dynamics across cryptocurrency categories. A data summary is presented in Table LABEL:tab:datacol.
@@ 4.2.1 Price Return Coherence with Sentiment
Figure 5: Rolling window for Bitcoin price in Long-term Coherence
DOGE: Figures 66(a) and 66(b) illustrate the coherence between Dogecoin price returns and sentiment trends on Tiktok and Twitter, respectively.
(a) Dogecoin price returns and TikTok sentiment coherence
(b) Dogecoin price returns and Twitter sentiment coherence
Figure 6: Rolling window for Dogecoin price in Short-term Coherence
USDT: Figures 77(a) and 77(b) illustrate the coherence between USDT price returns and sentiment trends on Tiktok and Twitter, respectively.
(a) USDT price returns and TikTok sentiment coherence
(b) USDT price returns and Twitter sentiment coherence
Figure 7: Rolling window for USDT price in Long-term Coherence
@@ 4.2.2 Volume Change Coherence with Sentiment
BTC: Figures 88(a) and 88(b) illustrate the coherence between Bitcoin volume changes and sentiment trends on Tiktok and Twitter, respectively.
@@
Figure 8: Rolling window for Bitcoin volume in Long-term Coherence
DOG: Figures 99(a) and 99(b) illustrate the coherence between Dogecoin volume changes and sentiment trends on Tiktok and Twitter, respectively.
(a) Dogecoin volume change and TikTok sentiment coherence
(b) Dogecoin volume change and Twitter sentiment coherence
Figure 9: Rolling window for Dogecoin volume in Short-term Coherence
USDT: Figures 1010(a) and 1010(b) illustrate the coherence between USDT volume changes and sentiment trends on Tiktok and Twitter, respectively.
(a) USDT volume change and TikTok sentiment coherence
(b) USDT volume change and Twitter sentiment coherence
Figure 10: Rolling window for USDT volume in Long-term Coherence
@@ 4.3.1 Unit root test
Table 2: Results of Unit Root for Price and Volume
Variable ADF Statistics Jarque-Bera
Price
BTCPRC 638.16*** -9.73**
ETHPRC 541.12*** -9.42**
SOLPRC 1476.42*** -9.32**
DOGPRC 6694.37*** -9.95**
XRPPRC 219593.15*** -9.60**
USDCPRC 4560786.4*** -14.41**
USDTPRC 50821.13*** -11.85**
Volume
BTCVOL 7099.2*** -9.74**
ETHVOL 172.26*** -13.67**
SOLVOL 1015519.61*** -13.12**
DOGVOL 98.6*** -15.35**
XRPVOL 309.11*** -14.77**
USDCVOL 22584.28*** -13.66**
USDTVOL 517.65*** -12.28**
Sentiment
Twitter Sentiment 129.0381*** -6.8333**
TikTok Sentiment 15.1845*** -8.0815**
@@ 4.3.2 Dynamic Connectedness and Pairwise Network Analysis
The network graphs in Figs. 11 and 12 further illustrate these dynamics. BTC, ETH, and SOL are the strongest transmitters of price return shocks, while XRP and DOGE act as net receivers. For volume changes, DOGE is the primary transmitter, influencing SOL and sentiment indices. Sentiment indices show limited outward influence but are highly sensitive to external shocks, consistent with Ji et al. [39]. The ETH-to-XRP connection observed in the network graph suggests evolving market relationships, aligning with Koutmos [82], who highlighted strong interdependencies among cryptocurrencies.
@@
Our results partially align with Li et al. [83], who noted that smaller-market-cap cryptocurrencies often transmit spillovers to larger ones. XRP follows this pattern, while DOGE deviates, emerging as the largest receiver of price return shocks but a key transmitter of volume changes. Fig. 12 illustrates DOGE’s impact on SOL volume and sentiment indices, while ETH volume also acts as a key transmitter to TikTok sentiment. These findings align with Balcilar et al. [84], who emphasized the critical role of volume in shaping market dynamics. The results highlight the interconnected nature of cryptocurrency markets and the importance of considering both price and volume dynamics in understanding market relationships.
Figure 11: Network graph for connectedness among sentiment and prices
Figure 12: Network graph for connectedness among sentiment and volumes
@@ 5.1 Case Study
The figures illustrate predictions for Dogecoin prices using TikTok (Figure 13) and Twitter (Figure 14) sentiment score as input signals. Both sets of predictions indicate that sentiments from TikTok and Twitter can follow Dogecoin’s price trends. However, significant differences emerge during periods of rapid price increases, such as in mid-October 2022. These variations highlight the importance of platform-specific sentiments in reflecting market behavior. This suggests that TikTok, with its younger and trend-sensitive user base, may be more effective in capturing speculative price surges. For traders, the results imply that incorporating TikTok sentiment could be useful for monitoring short-term price fluctuations. For instance, speculative investors might rely more heavily on TikTok sentiment to identify quick price movements, while institutional or long-term investors may benefit from using both TikTok and Twitter sentiments to gain a broader understanding of market dynamics.
Figure 13: Prediction DOG price with Tiktok
Figure 14: Prediction on DOG price with Twitter

### v2

@@ Abstract
[inst2]organization=School of Information Systems, Queensland University of Technology, city=Brisbane, state=Queensland, country=Australia
@@
We introduce a framework for analyzing TikTok video-based sentiment compared to Twitter text-based sentiment in cryptocurrency sentiment analysis.
We find that TikTok sentiment particularly influences short-term speculative trends, while Twitter sentiment aligns better with long-term market dynamics.
@@ 3.1 Data Collection
Account Name crypto.lucho CryptoMasun CryptoProjects thecryptonetwork cryptosanti
Number of Followers 2.7M 1.4M 1.3M 1.0M 914K
Account Name cryptorobles suppomancrypto Virtual Bacon thecryptohippie Cryptowendyo
Number of Followers 886K 344K 338K 288K 277K
@@
For each group, one representative coin was chosen to conduct dependency analysis between sentiment signals and forecast modeling, providing insights into the differing dynamics across cryptocurrency categories. A data summary is presented in Table 2.
Table 2: Data Summary for TikTok and Twitter
Platform Period Number of data Average likes Average comments Average shares
TikTok 2021.11–2024.03 15,348 1,350 210 90
Twitter 2021.11–2024.03 519,208 240 35 15
@@ 4.2.1 Price Return Coherence with Sentiment
DOGE: Figures 55(c) and 55(d) illustrate the coherence between Dogecoin price returns and sentiment trends on Tiktok and Twitter, respectively.
USDT: 55(e) and 55(f) illustrate the coherence between USDT price returns and sentiment trends on Tiktok and Twitter, respectively.
@@
(c) Dogecoin price returns and TikTok sentiment coherence
(d) Dogecoin price returns and Twitter sentiment coherence
(e) USDT price returns and TikTok sentiment coherence
(f) USDT price returns and Twitter sentiment coherence
Figure 5: Rolling window coherence between cryptocurrency price returns and social media sentiment: (a,b) Bitcoin (long-term), (c,d) Dogecoin (short-term), (e,f) USDT (long-term).
@@ 4.2.2 Volume Change Coherence with Sentiment
BTC: Figures 66(a) and 66(b) illustrate the coherence between Bitcoin volume changes and sentiment trends on Tiktok and Twitter, respectively.
@@
DOG: Figures 66(c) and 66(d) illustrate the coherence between Dogecoin volume changes and sentiment trends on Tiktok and Twitter, respectively.
USDT: Figures 66(e) and 66(f) illustrate the coherence between USDT volume changes and sentiment trends on Tiktok and Twitter, respectively.
@@
(c) Dogecoin volume change and TikTok sentiment coherence
(d) Dogecoin volume change and Twitter sentiment coherence
(e) USDT volume change and TikTok sentiment coherence
(f) USDT volume change and Twitter sentiment coherence
Figure 6: Rolling window coherence between cryptocurrency volume changes and social media sentiment: (a,b) Bitcoin (long-term), (c,d) Dogecoin (short-term), (e,f) USDT (long-term).
@@ 4.3.1 Unit root test
Table 3: Results of Unit Root Tests for Price, Volume, and Sentiment
Test BTCPRC ETHPRC SOLPRC DOGPRC XRPPRC USDCPRC USDTPRC
ADF Statistics 638.16*** 541.12*** 1476.42*** 6694.37*** 219593.15*** 4560786.4*** 50821.13***
Jarque–Bera -9.73** -9.42** -9.32** -9.95** -9.60** -14.41** -11.85**
Test BTCVOL ETHVOL SOLVOL DOGVOL XRPVOL USDCVOL USDTVOL
ADF Statistics 7099.2*** 172.26*** 1015519.61*** 98.6*** 309.11*** 22584.28*** 517.65***
Jarque–Bera -9.74** -13.67** -13.12** -15.35** -14.77** -13.66** -12.28**
Test Twitter Sent. TikTok Sent.
ADF Statistics 129.04*** 15.18***
Jarque–Bera -6.83** -8.08**
@@ 4.3.2 Dynamic Connectedness and Pairwise Network Analysis
The network graphs in Figs. 7(a) and 7(b) further illustrate these dynamics. BTC, ETH, and SOL are the strongest transmitters of price return shocks, while XRP and DOGE act as net receivers. For volume changes, DOGE is the primary transmitter, influencing SOL and sentiment indices. Sentiment indices show limited outward influence but are highly sensitive to external shocks, consistent with Ji et al. [39]. The ETH-to-XRP connection observed in the network graph suggests evolving market relationships, aligning with Koutmos [82], who highlighted strong interdependencies among cryptocurrencies.
@@
Our results partially align with Li et al. [83], who noted that smaller-market-cap cryptocurrencies often transmit spillovers to larger ones. XRP follows this pattern, while DOGE deviates, emerging as the largest receiver of price return shocks but a key transmitter of volume changes. Fig. 7(b) illustrates DOGE’s impact on SOL volume and sentiment indices, while ETH volume also acts as a key transmitter to TikTok sentiment. These findings align with Balcilar et al. [84], who emphasized the critical role of volume in shaping market dynamics. The results highlight the interconnected nature of cryptocurrency markets and the importance of considering both price and volume dynamics in understanding market relationships.
(a) Connectedness among sentiment and prices
(b) Connectedness among sentiment and volumes
Figure 7: Network graphs for connectedness among sentiment, prices, and volumes.
@@ 5.1 Case Study
The figures illustrate predictions for Dogecoin prices using TikTok (Figure 8(a)) and Twitter (Figure 8(b)) sentiment score as input signals. Both sets of predictions indicate that sentiments from TikTok and Twitter can follow Dogecoin’s price trends. However, significant differences emerge during periods of rapid price increases, such as in mid-October 2022. These variations highlight the importance of platform-specific sentiments in reflecting market behavior. This suggests that TikTok, with its younger and trend-sensitive user base, may be more effective in capturing speculative price surges. For traders, the results imply that incorporating TikTok sentiment could be useful for monitoring short-term price fluctuations. For instance, speculative investors might rely more heavily on TikTok sentiment to identify quick price movements, while institutional or long-term investors may benefit from using both TikTok and Twitter sentiments to gain a broader understanding of market dynamics.
(a) Prediction on DOG price with TikTok
(b) Prediction on DOG price with Twitter
Figure 8: Predictions of DOG price with sentiment signals from TikTok and Twitter.